# Magnetocaloric effect for a $Q$-clock type system

### v1

@@ Abstract
In this work, we study the magnetocaloric effect applied to a magnetic working substance corresponding to a square lattice of spins with $Q$ possible orientations known as the “$Q$-state clock model” where for $Q\geq 5$, the systems present the famous Berezinskii–Kosterlitz–Thouless phase (BKT). Thermodynamic quantities are obtained in exact form for a small lattice size of $L\times L$ with $L=3$ and by the mean-field approximation and Monte Carlo simulations for $Q$ pairs between 2 and 8 with $L=3,8,16,32$ with free boundary conditions, and magnetic fields varying between $B=0$ and $1$ in natural units of the system. By obtaining the entropy, it is possible to quantify the caloric effect through an isothermal process in which the external magnetic field on the spin system is varied. In particular, we find the values of $Q$ that maximize the effect depending on the lattice size and the magnetic phase transitions related to maximizing the caloric phenomena. These indicate that in a small lattice (up to $\sim 7\times 7$), when $Q\geq 5$, the transition that maximizes the effect is related to ferromagnetic to BKT type. In contrast, transitioning from BKT to paramagnetic type increases the system’s caloric response when we work with a larger lattice size.
@@ I Introduction
The caloric phenomena constitute a fundamental axis of study in the physics of materials to find replacement candidates for the toxic gases used in conventional refrigeration systems [1, 2]. To this end, considerable temperature variations have been achieved for different types of materials after an adiabatic variation of the value of the applied hydrostatic pressure (barocaloric effect) [3, 4, 5, 6], mechanical stress (elastocaloric effect) [7, 8, 9, 10, 11], electric field (electrocaloric effect) [12, 13, 14, 15] and external magnetic field (magneto-caloric effect, MCE) [16, 17, 18, 19]. Although each phenomenon’s nature and physical interpretation are different, they comply with a basic concept that characterizes them: the aim is to maximize the variation of entropy in the face of changes in the control parameter that governs the thermodynamic process that describes the phenomenon [20, 21, 22, 23, 24, 25, 26]. In particular, the systems that present phase transitions maximize entropy variation near the transition [27, 28, 29, 30]. One of many corresponds to magnetic materials, characterized by a determined magnetic order or ground state. Remarkably, such an order depends on the energy contributions in the material [31, 32, 33]. Within these transitions, we will focus on those magnetic systems that present as the temperature increases and thus change their magnetic phase. An interesting system to explore corresponds to the $Q$-state clock model that constitutes the discrete version of the famous 2D XY model [34, 35, 36, 37], which is probably the most extensively studied example showing the Berezinskii–Kosterlitz–Thouless (BKT) transition in the presence of a frustrated quenched disordered phase [38, 39, 40]. The $Q-$state clock model is one of many magnetic models to mimic the thermodynamics of some materials, and it can be viewed as a classical Heisenberg spins model with very strong planar anisotropy [41, 42, 43].
One way to characterize the phase transitions of the $Q$-clock state model is through the maxima obtained in the specific heat as a function of temperature. Each location of a maximum of the specific heat on the temperature axis will represent a value for a so-called critical temperature. It has been shown [44, 45, 46, 47, 48, 49] (in the absence of an external magnetic field) that for the $Q$-clock state model, values $Q\geq 5$ (where $Q$ represents the number of possible orientations that the spins can take), the specific heat presents two maxima. The first maximum corresponds to transitioning from a ferromagnetic phase (FP) to a BKT phase. In contrast, the second maximum transitions from BKT to a paramagnetic disordered phase (PP) [36].
In this sense, this model, which presents these two phase transitions mentioned above, is very interesting to exploit in studying caloric phenomena. In addition, applying this spin model to this effect raises several questions: which value of $Q$ is the optimum, i.e., maximizes the thermodynamic response? How does the estimation of the magnitude of the effect through mean-field approximation compare to the exact calculations and Monte Carlo simulations? Is it possible to know which magnetic phases maximize the effect? Is there convergence in the results when we increase the size of the lattice?
One of the technical difficulties with this model is the scaling of the accessible states, which scales as $\sim Q^{L^{2}}$ where $L\times L$ represents the size of the lattice you work. This means that when solving thermodynamics exactly via a canonical partition function, the computational costs result in long computation times. With this limitation, exact model computations can only be applied to lattices of small size. That is why two relevant mechanisms are employed to see what happens with larger lattices: Mean field approximation and Monte Carlo simulations.
In this work, we study the MCE in the $Q$-clock model, in which the thermodynamic quantities of the model are obtained exactly for small lattices and using a mean-field approximation and Monte Carlo simulations for larger lattice sizes. The $Q$ values studied correspond to the even values $Q=2,4,6$ and $8$. The exact calculation for a small lattice of $L\times L$ with $L=3$ is presented to review the MCE concerning the magnetic phase and to quantify the caloric effect per spin of small-size lattices. In completeness, the results are analyzed using the mean-field approximation of this lattice to verify if the estimates in the orders of magnitude of the effect are consistent with this approximation. Finally, our analysis is extended to $16\times 16$ and $32\times 32$ lattices employing a mean-field and Monte Carlo simulations.
Our work is structured as follows: Section II presents the $Q$-clock model, Section III revisits the mean-field approximation, Section IV analyzes the thermodynamic quantities involved in the calculations, Section V gives an analysis of the caloric phenomena, Section VI shows the results, and finally, Section VII presents our work’s conclusions.
@@ II.1 $Q$-state clock model
Figure 1: (a) $6\times 6$ spin lattice and (b) $Q$ clock model for $Q=2,4,6$ and $8$.
The working substance under study corresponds to the $Q$-states clock model on a two-dimensional (2D) square lattice of dimensions $L\times L=N$, where the local magnetic moment or “spin” $S_{i}$ at site $i$ can point in any of $Q$ directions in a given plane. $S_{i}$ is then a 2D vector, i.e. $S_{i}=(\cos(\frac{2\pi}{q}k),\sin(\frac{2\pi}{q}k))$, where $k=0,1,...Q-1$, with equal probability. The magnitude of $S_{i}$ is chosen as the unity.
@@
$\mathcal{H}=-\sum_{\langle i,j\rangle}\mathcal{J}(\vec{S}_{i}\cdot\vec{S}_{j})-\sum_{i}\vec{B}\cdot\vec{S}_{i}\,,$ (1)
where $\mathcal{J}>0$ is the ferromagnetic exchange interaction to nearest neighbors; the sum runs over all pairs of nearest neighbors $(i,j)$, which is indicated by the symbol $\langle i,j\rangle$ under the summation symbol. $\vec{B}$ is an external field applied along one direction in the plane. In this work, we used arbitrary units choosing $\mathcal{J}=1$, making all the calculated quantities for the exchange energy constant. From this Hamiltonian model, we must extract the accessible states of the system to obtain the entropy of the system.
To quantify the caloric effect, it is necessary to obtain the entropy from the thermodynamic study of the model. We will first analyze two cases where this calculation comes from a canonical partition function. The first corresponds to an exact calculation of all the accessible microstates of the system for a $3\times 3$ lattice for $Q=2,4,6$ and $8$. For the same values of $Q$ studied, the mean-field theory will be employed for lattices of size up to $32\times 32$. On the other hand, Monte Carlo simulations offer us the possibility to obtain thermal averages with a sample of the configurations and to define an entropy via integration of the specific heat.
In the next section, we will review the mean-field approximation to emphasize the calculation of the approximate partition function and the definition taken from there for the entropy coming from the Helmholtz free energy temperature derivative, which also applies to the exact case in small lattices.
III Mean Field Approximation

### v2

@@ Abstract
In this work, we study the magnetocaloric effect (MCE) in a working substance corresponding to a square lattice of spins with $Q$ possible orientations, known as the “$Q$-state clock model”. When the $Q$-state clock model has $Q\geq 5$ possible configurations, it presents the famous Berezinskii–Kosterlitz–Thouless (BKT) phase associated with vortices states. We calculate thermodynamic quantities using Monte Carlo simulations for even $Q$ numbers, ranging from $Q=2$ to $Q=8$ spin orientations per site in a lattice. We use lattices of different sizes with $L\times L=8^{2},16^{2},32^{2},64^{2},\text{and}\ 128^{2}$ sites, considering free boundary conditions and an external magnetic field varying between $B=0$ and $B=1$ in natural units of the system. By obtaining the entropy, it is possible to quantify the MCE through an isothermal process in which the external magnetic field on the spin system is varied. In particular, we find the values of $Q$ that maximize the MCE depending on the lattice size and the magnetic phase transitions linked with the process. Given the broader relevance of the $Q$-state clock model in areas such as percolation theory, neural networks, and biological systems, where multi-state interactions are essential, our study provides a robust framework in applied quantum mechanics, statistical mechanics and related fields.
@@ I Introduction
Caloric phenomena form a foundational aspect of material physics, crucial in identifying viable substitutes for the toxic gases currently employed in conventional refrigeration systems [1, 2]. Significant temperature variations driven by caloric processes have been observed in different materials following adiabatic changes in applied hydrostatic pressure (barocaloric effect) [3, 4, 5, 6], mechanical stress (elastocaloric effect) [7, 8, 9, 10, 11], electric field (electrocaloric effect) [12, 13, 14, 15], and external magnetic field (magnetocaloric effect, MCE) [16, 17, 18, 19]. Although the underlying nature and physical interpretation of each phenomenon differ, they share a common objective: maximizing entropy changes in response to variations of the control parameter governing the thermodynamic process [20, 21, 22, 23, 24, 25, 26].
Systems exhibiting phase transitions, in particular, achieve maximum entropy changes near where these transitions are occurring [27, 28, 29, 30]. Among these, magnetic materials stand out due to their distinct magnetic order or ground state, which is heavily influenced by the energy contributions within the material [31, 32, 33]. Here we concentrate on magnetic systems that undergo phase changes with increasing temperature, specifically those that alter their magnetic phase. A notable model for exploration is the $Q$-state clock model, a discrete variant of the renowned 2D XY model [34, 35, 36, 37], widely studied for illustrating the Berezinskii–Kosterlitz–Thouless (BKT) transition in frustrated quenched disordered phases [38, 39, 40]. The $Q$-state clock model serves as a classical Heisenberg spin model with pronounced planar anisotropy, useful in replicating material thermodynamics [41, 42, 43].
Phase transitions in the $Q$-clock state model can be identified via the maxima observed in the specific heat as a function of temperature. Each maximum signifies there is a ”critical temperature.” In scenarios devoid of an external magnetic field, studies have demonstrated [44, 45, 46, 47, 48, 49] that for $Q\geq 5$, the specific heat exhibits two maxima. The initial maximum pertains to the transition from a ferromagnetic phase (FP) to a BKT phase, followed by a transition from BKT to a paramagnetic disordered phase (PP) [36].
Due to these dual phase transitions, the $Q$-clock state model is particularly intriguing for caloric phenomenon studies. This model prompts several critical questions: What is the optimal value of $Q$ that maximizes entropy variation and, consequently, the MCE? Which magnetic phases enhance the effect most effectively? Are the results consistent as the lattice size increases?
A major technical challenge with this model revolves around scaling accessible states, which scales approximately as $\sim Q^{L^{2}}$, where $N=L\times L$ defines the lattice size with $N$ number of sites. Consequently, exact computations using the canonical partition function entail substantial expenses, thus restricting exact computations to smaller lattices. Mean Field approximation and Monte Carlo simulations allow us to address large lattice size challenges. Monte Carlo simulations are often preferred for their accuracy in representing systems with intricate interactions and fluctuations, which the mean-field approach may overlook due to its averaging assumptions. Consequently, mean-field theory may not aptly describe short-range interactions like those in the systems examined herein.
Experimental validation of the $Q$-state clock model is complex due to its theoretical nature. Nonetheless, experimental scenarios reflecting behaviors predicted by this model, especially considering phase transitions, exist, notably in magnetic thin films and two-dimensional magnets [35], liquid crystals [50], and cold atom systems [51].
This work investigates the MCE in the $Q$-clock model, utilizing Monte Carlo simulations to derive the model’s thermodynamic properties for large lattice sizes. The analyzed $Q$ values are even numbers ($Q=2,4,6,8$), facilitating exploration into how the spin orientation given by $Q$ numbers influences the magnetocaloric effect across variations of thermal and magnetic conditions. We examine the magnetocaloric effect’s reliance on these $Q$ values, focusing on entropy, specific heat, internal energy, and magnetization responses near phase transitions. This study elucidates optimal spin configurations to maximize the MCE in discrete multi-state systems.
Our work is structured as follows: Section II introduces the $Q$-state clock model, detailing its Hamiltonian and the relevant thermodynamic quantities. Section III describes the Monte Carlo simulation techniques used to derive thermodynamic properties, including detailed explanations of lattice sizes, boundary conditions, and sampling methodologies. Section IV examines the calculated thermodynamic properties, such as specific heat, internal energy, and entropy, for various values of $Q$. Section V explores the magnetocaloric effect and outlines the methodology for its calculation. Section VI presents our findings, highlighting the optimal conditions and $Q$ values that maximize the effect. Finally, Section VII summarizes the key findings of our research.
@@ II.1 $Q$-state clock model
Figure 1: (a) Schematic representation for a square lattice of size $6\times 6$ with spin orientations corresponding to $Q=8$. (b) $Q$-clock model for $Q=2,4,6$ and $8$ states. Orange dots indicate sites, and blue arrows the possible spin orientation at each site with spin vector $\vec{S_{i}}$.
The system under study corresponds to the $Q$-states clock model on a two-dimensional (2D) square lattice of dimensions $L\times L=N$ (with $N$ the total number of sites in the lattice). The local magnetic moment or “spin” $\vec{S}_{i}$ at site $i$ can point in any of $Q$ directions in a given plane, see Fig. 1. $\vec{S}_{i}$ is then a 2D vector, i.e. $\vec{S}_{i}=(\cos(\frac{2\pi}{Q}k),\sin(\frac{2\pi}{Q}k))$, where $k=0,1,...Q-1$, with equal probability. The magnitude of $\vec{S}_{i}$ is chosen as the unity.
@@
$\mathcal{H}=-\sum_{\langle i,j\rangle}\mathcal{J}\vec{S}_{i}\cdot\vec{S}_{j}-\sum_{i}\vec{B}\cdot\vec{S}_{i}\,,$ (1)
where $\mathcal{J}>0$ is the ferromagnetic exchange interaction to nearest neighbors; the sum runs over all pairs of nearest neighbors $\langle i,j\rangle$. $\vec{B}$ is an external magnetic field applied along one of the directions of the plane. We use $\mathcal{J}=1$ in our calculations, thus all quantities are delivered in exchange units.
To quantify the caloric effect, the entropy from the thermodynamics of the model must be obtained. To do that, we will analyze Monte Carlo simulations, which offer the possibility of obtaining thermal averages within a sample of the possible configurations of the system. The following section will detail the method employed, including the algorithm used in our computational calculations.
III Monte Carlo Simulations
The work presented herein will focus on numerical computations based on Monte Carlo (MC) simulations. A square lattice of size $L\times L$ is selected, with free boundary conditions imposed. A site is randomly chosen for visitation, and the energy cost, $\Delta$, associated with rotating the corresponding spin among $Q$ possible states is calculated. If the energy is diminished, the change in orientation is accepted; otherwise, the spin rotation is accepted only when $\exp(-\Delta/T)\leq r$, where $r$ is a freshly generated random number uniformly distributed in the range [0,1]. This procedure conforms to the conventional Metropolis algorithm. A Monte Carlo step (MCS) is achieved after $N=L\times L$ spin-rotation attempts.
For each lattice size and $Q$ value, a sequence of temperatures has been established within the range [0.02, 4] in increments of 0.02 for each temperature. A total of $5\tau$ MCSs are performed: the initial $\tau$ MCSs are utilized to equilibrate the system at a fixed temperature $T$. In contrast, the subsequent $4\tau$ MCSs are employed to measure observables every 20 MCSs, achieving a cumulative total of $200.0002\times 10^{5}=$ 200.000 measurements. From now on, $\tau=10^{6}$, unless otherwise specified for the remainder of this paper, this choice of $\tau$ yields stable results and corresponds with the analytical expressions obtained in previous studies for smaller lattices.
III.1 Thermal averages
The magnetization per site $\vec{M}$ is given by
$\vec{M}=\frac{1}{N}\sum_{i=1}^{N}\vec{S}_{i},$ (2)
where $\vec{S}_{i}$ is the spin at site $i$ at a given time, $t$, and $N=L\times L$. In this case $\vec{M}$ is a vector of two components $\vec{M}=(M_{x},M_{y})$. Normally, this vector’s magnitude or absolute value is calculated as $|\vec{M}|=\sqrt{M_{x}^{2}+M_{y}^{2}}$. Then, the thermal average of $|\vec{M}|$ is $<|\vec{M}|>$ given by
$<|\vec{M}|>=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\sqrt{M_{x}^{2}+M_{y}^{2}},$ (3)
with $N_{c}=2\times 10^{5}$ the number of configurations used to perform thermal averages.
Energy is the main quantity the Monte Carlo method uses to reach thermal equilibrium. Therefore, after $\tau$ MCSs the internal energy $U$ can be obtained by averaging the $N_{c}=2\times 10^{5}$ values for $E_{k}$, where $k$ runs over the accepted configurations after the Metropolis algorithm, namely
$U=<E>=\frac{1}{N_{c}}\sum_{k=1}^{N_{c}}E_{k},$ (4)
where every spin configuration is separated from the next one by $20$ MCSs. The energy per site is then $U/N$.
The specific heat is then calculated as proportional to the fluctuations of the energy as follows (we use $k_{B}=1$ for simplicity):
$C=\frac{\langle E^{2}\rangle-\langle E\rangle^{2}}{T^{2}},$ (5)
$C=\frac{1}{T^{2}}\left[\left(\frac{1}{N_{c}}\sum_{k=1}^{N_{c}}E_{k}^{2}\right)-\left(\frac{1}{N_{c}}\sum_{k=1}^{N_{c}}E_{k}\right)^{2}\right].$ (6)
The entropy $S$ can be calculated by numerical integration of the specific heat over $T$ as
[…]
Figure 2: (a) Normalized internal energy $U/N$ as a function of temperature for $Q=2$ and lattice sizes $8\times 8$, $16\times 16$, $32\times 32$, $64\times 64$ and $128\times 128$ for an external magnetic field of (a) $B=0$, (b) $B=1$.
III.2 Selection of the lattice size
Due to the scaling of the system’s accessible states ($\sim Q^{L^{2}}$ ), it is necessary to set a criterion for a representative lattice size in the thermodynamic limit, where the results of the caloric studies are valid and closer to reality. For this purpose, we have analyzed the convergence in the internal energy $U$ for different values of the lattice size from our Monte Carlo simulations. For this purpose, we have decided to explore the internal energy of the system (normalized to the number of sites) as a function of $T$ for $Q=2$ and different lattice sizes, corresponding to $8\times 8$, $16\times 16$, $32\times 32$, $64\times 64$ and $128\times 128$, and for magnetic field values of $B=0$ and the largest field that we will consider $B=1$. This is observed in Fig. 2(a)-(b), where it can be seen that for a lattice size of $128\times 128$, the variation between their internal energy compared to the $64\times 64$ lattice is extremely small for both $B$ values. Therefore, a larger lattice, for example, $256\times 256$, would no longer show substantial differences in internal energy as in comparison to a $128\times 128$ lattice, within the precision of Fig. 2, which is enough for the scope of this paper. That is why our numerical thermodynamic limit will correspond in the following analysis to a lattice size of $128\times 128$.
IV Analysis of Thermodynamic Quantities
Figure 3: Normalized specific heat $C/N$ as a function of temperature for even $Q$ values between $Q=2$ and $Q=8$ and a $128\times 128$ lattice. (a) $B=0$, (b) $B=1$. The shift of the peaks to the right at $B=1$ is clearly seen for all $Q$.
We will begin by reviewing the thermodynamic results of the model, focusing first on the analysis of the specific heat for a $128\times 128$ lattice. This thermodynamic quantity for a field $B=0$ and field $B=1$ is shown in Fig. 3(a) and Fig. 3(b), respectively, for $Q=2,4,6,8$. Here, we observe two peaks for $Q\geq 6$, indicative of a double-phase transition. It is, therefore, the BKT phase that appears in these cases. It is also noticeable that the critical temperature of each transition increases as $B$ increases, and their associated peaks are less pronounced, as we can observe when comparing Fig. 3 panel (a) with panel (b). This is because the external magnetic field favors ordered phases (FM and BKT) over disordered ones.
Following with the $128\times 128$ lattice, its internal energy as a function of $T$ is shown in Fig. 4(a), where we can observe that lower values are obtained when $B=1$ (blue curves) than when $B=0$ (purple curves). This is due to the Zeeman term present in the Hamiltonian given by Eq. (1), which decreases the ground state energy compared to the case in the absence of a magnetic field. Figure 4(b) shows the magnetization $M$ of the system for $Q=2,4,6,$ and $8$ for $B=1$ (blue curves) and $B=0$ (purple curves). First, we notice that all systems start saturated (in a ferromagnetic state), and then $M$ decreases as $T$ increases. The cases with $B=1$ show higher magnetization than those with $B=0$ as $T$ increases for all $Q$ displayed. It is observed that the change in magnetization as a function of temperature happens much faster for larger $Q$ values. This is a direct consequence of the degrees of freedom of the system. For example, we can think of $Q=2$, where two possible orientations for the spin are up and down. In terms of energy, generating a flip of such a configuration is much more difficult. Therefore, the magnetization will be higher for low $Q$ than for higher values of $Q$ at the same temperature.
Figure 4: Normalized (a) internal energy, (b) magnetization, and (c) entropy as a function of temperature for even values of $Q$ between $Q=2$ and $Q=8$ and a $128\times 128$ lattice. $B=0$ (purple curves), $B=1$ (blue curves).
Next, we analyze the entropy in Fig. 4(c). We observe that for a given $Q$, the entropy is lower when $B=1$ (blue curves) than for the case of $B=0$ (purple curves). This is consistent with the interpretation that the magnetic field tends to order the spin system. Also, it is noticeable that for more spin degrees of freedom (as $Q$ increases), the entropy is higher as a function of temperature, which is consistent with the more significant number of accessible states the system acquires as $Q$ increases. These curves show that the entropy fulfills $\lim_{T\rightarrow 0}S(T,B\neq 0)=0$ because, with even a very low field value, the system at low temperature tends to have only one preferred state due to the presence of the external field applied to the system. It is important to remember that in the case of a null field, the entropy at low temperature has the value of $\lim_{T\rightarrow 0}S(T,0)=\ln Q$. It should be noted that the entropy in Fig. 4 is normalized to the total number of spins in the lattice, i.e., $S(T,B)/N$. This implies that the quantity $S_{0}/N$ is very small when the field tends to zero (and strictly zero with $B\neq 0$). In the case of lattice size $128\times 128$ displayed in Fig. 4, this number would be given by $\ln(Q)/16384$, in which the most significant value would be provided for $Q=8$, offering a correction of the order of $10^{-4}$.
V Magnetocaloric Phenomena
To analyze the magnetocaloric effect, we start by treating the total entropy of a material as the sum of three entropies: electronic $S_{e}$, magnetic $S_{m}$, and lattice $S_{l}$.
$S(T,B)=S_{e}(T,B)+S_{m}(T,B)+S_{l}(T,B).$ (8)
It is essential to mention that the latter equation presumes the separation of the orbital, magnetic, and lattice degrees of freedom. This assumption does not always hold, as evidenced by systems exhibiting Jahn-Teller interactions or magnetoelastic coupling, where these degrees of freedom are intertwined [52, 53, 54, 55].
Magnetic entropy $(S_{m})$ strongly depends on the magnetic field. Conversely, it is typically observed that many materials’ electronic and lattice entropies are mainly independent of the magnetic field. However, this is not universally applicable; notably, in low-temperature regimes (approximately below 10 K), some studies suggest that the electronic entropy may exhibit nonlinear dependence on the magnetic field [56].
In this study, we consider a more conventional system where the magnetic field does not influence the lattice and electronic entropies, allowing us to reformulate Eq. (8) as
$S(T,B)\simeq S_{e}(T)+S_{m}(T,B)+S_{l}(T).$ (9)
The quantification of the caloric phenomenon is associated with a thermodynamic refrigerator cycle, where we can take two paths to quantify the effect: i) an adiabatic trajectory and ii) an isothermal trajectory. The temperature variation that suffers the systems along the isentropic process is $\Delta T_{ad}$, and is given by
[…]
where we use that $C_{B}=\left(\frac{\partial U}{\partial T}\right)_{B}=T\left(\frac{\partial S}{\partial T}\right)_{B}$ corresponds to specific heat at constant $B$.
In the case of quantifying the effect employing an isothermal trajectory, we obtain an entropy variation at constant temperature, $\Delta S$, given by
[…]
If we look at Eq. (10) and Eq. (11), we can find a relationship between these quantities. It is found that $-\Delta S\propto\Delta T_{ad}$. Consequently, it is essential to note that when one has a case in which $-\Delta S>0$, let’s call this kind of response of direct type. The system will heat up, while when the response is $-\Delta S<0$ type, we call this inverse type response, and consequently, the system will cool down. Therefore, we expect in a direct response a $\Delta T_{ad}>0$, and for the case of an inverse response, we expect a $\Delta T_{ad}<0$ for the final result of the caloric phenomena.
For the case of the entropy variation at constant temperature $T$, the magnetocaloric expression can be given by the difference of the entropy at the initial and final point of the process as
$-\Delta S\simeq-\Delta S_{m}\simeq S_{m}(T,B_{0})-S_{m}(T,B),$ (12)
where the contributions of $S_{l}$ and $S_{e}$ seen in Eq. 9 vanish due to their only temperature dependence, and consequently, in an isothermal process (as the one we analyze here), they do not have associated variations as $B$ changes. Equation 12 generates a graph of $-\Delta S$ as a function of $T$ for a given $B_{0}$ and final $B$ that quantifies the magnetocaloric phenomena.
For temperature and magnetic field ranges we use in our calculations, we fix the zero value of temperature at $T=0.01$, while the smallest field we work is $B=0$. On the other hand, the maximum temperature corresponds to $T=4.02$, and magnetic field of $B=1$, all in $\mathcal{J}$ units.
VI Results
VI.1 Direct or indirect caloric response?
Figure 5: Normalized entropy difference $-\Delta S=S(T,B_{0}=0)-S(T,B)$ as a function of temperature for lattices sizes from $8\times 8$ to $128\times 128$ sites, and $B=0.2$ (lower curves) and $B=1.0$ (upper curves). (a) $Q=2$, (b) $Q=8$.
To understand the caloric response of the system, let us start the analysis with a simple example for $Q=2$, top left panel in Fig. 1(b), and lattices increasing in size from $8\times 8$ to $128\times 128$. The entropy difference that quantifies the MCE in Eq. 12 is given by $-\Delta S=S(T,0)-S(T,B)$, where we have selected $B_{0}=0$. This entropy difference is plotted for different final magnetic fields of $B=0.2$ and $B=1.0$ as a function of the temperature $T$ the process is taken, as shown in Fig. 5(a). We note that the response of the effect is positive for both presented cases of different $B$, which indicates a direct type behavior for the MCE. It is also observed that for a lattice of $128\times 128$, and below $T\sim 1.3$, the thermal response of the effect for $B=0.2$ and $B=1.0$ is the same. This latter also occurs for smaller lattices but for lower temperatures than $1.3$.
As temperature increases after $T\sim 1.3$, it is observed that the case for $B=1$ is notably different from the one with $B=0.2$, as the peaks for $B=1$ are larger and shifted to higher $T$ for all lattice sizes. The explanation of this phenomenon is relatively simple. The system prefers to be in a ferromagnetic configuration (i.e., in a single configuration state) in an external magnetic field up to reach a maximum in $-\Delta S$. To remove the spin ordering from this state requires a considerable increase in temperature in terms of energy. This is reflected in the entropy for $Q=2$ and $128\times 128$ lattice for $B=0$ and $B=1$, where both entropies are equal up to $T\sim 1.3$, and then they separate as $T$ increases, see two lower curves in Fig. 4(c).
Using the same criteria previously discussed, we will obtain similar behavior associated with the sign of $-\Delta S$ for different $Q$ values. That is to say, we will obtain a response of the direct type $-\Delta S>0$ for $Q=8$ even independent of the lattice size, as can be appreciated in Fig. 5(b). This is because the entropy without an external magnetic field is always greater than the entropy for any non-zero magnetic field value for any value of $Q$, as we can see for the lattice of $128\times 128$ sites in Fig. 4(c).
In addition, we observe by comparing both panels (a) and (b) of Fig. 5, that the maxima for $-\Delta S/N$ with $32\times 32$ and $128\times 128$ lattice sites do not present significant variations in magnitude neither the $T$ value where they are located in the horizontal axis. This statement is even more noticeable in the case of $Q=8$ in panel (b), indicating that as the lattice size and $Q$ increases, the entropy changes curves tend to be similar for the $T$ range and $B$ values we are considering in these calculations. It is also important to emphasize that an inverse effect $-\Delta S<0$ can be obtained in an isothermal process, starting from a higher magnetic field and going to a lower magnetic field. Still, it is much more natural to have a system without a magnetic field and activate an external field on the spin system afterward.
For the larger $Q$ value we are considering $Q=8$ in Fig. 5 panel (b), it is observed that the MCE is more significant for the second transition in all presented lattice sizes, as seen by the larger maxima in the temperature range from $T=1$ to $T=1.5$. This is true for large lattices constructed with Monte Carlo simulations and mean-field theory. However, using an exact formulation performed on the canonical ensemble, the opposite is true for small lattices (e.g., a $3\times 3$ lattice, see Fig. 7(b) for $Q=8$). Since we are focusing on the thermodynamic limit, we will not delve further into the effect for small lattices in the main text of our work. However, we have added a full Appendix A on exact and mean field calculations for small systems to complement this study.
VI.2 What $Q$ value has the better magnetocaloric response?
In order to answer which value of $Q$ maximizes the magnetocaloric effect, we have to consider two temperature regimes. Since in our model, temperature is in units of $\mathcal{J}$, it is convenient to think of two temperature regimes, the first one for $10<T<$ 1, and the second one for $T>1$.
From Fig. 6, we observe the magnetocaloric response for the $128\times 128$ lattice with even $Q$ values. Due to the double-phase transition in the model for $Q>5$, each $Q=6$ and $Q=8$ (circle symbols) display two distinct maxima in the caloric effect as a function of $T$. The first maximum (from left to right in $T$) is associated with a FP-BKT phase transition. The second maxima with a BKT-PP transition. For both $Q=6$ and $Q=8$, at temperatures $T<1$, the magnitudes of $-\Delta S$ are nearly the same, but the first maximum occurs for $Q=8$, reaching its peak at a lower temperature than $Q=6$. This trend aligns with the expected behavior of our model, where increasing the spin degrees of freedom $Q$ allows for similar caloric responses at progressively lower temperatures. As $Q$ increases, the temperature at which the FP-BKT transition occurs shifts leftward, approaching to zero in the theoretical limit as $Q$ tends to infinity. In the higher temperature region $T>1$, we observe that the caloric effect for $Q=8$ slightly surpasses that of $Q=6$, particularly near the BKT-PP transition. This suggests that, for $Q>5$, the BKT-PP transition becomes more pronounced with increasing spin degrees of freedom $Q$, amplifying the caloric response as $Q$ grows.
Figure 6: Normalized magnetocaloric effect for a $128\times 128$ lattice using Monte Carlo simulations for the case of $Q=2$ up to $Q=8$ and external $B=1$.
If we only focus on the maximum magnitude of the MCE and not its location in temperature, this is given for $Q=4$ for any lattice size. This is due to the sharp peak observed in the specific heat for $Q=4$ without an external magnetic field, as seen in Fig. 3(a). If we compare this peak with the one obtained when $B=1$ (Fig. 3(b)), the magnitude peak decreases almost by half, the most remarkable decrease of all the peaks observed for all the $Q$ studied. This will generate significant entropic differences and a higher associated caloric effect, as shown in Fig. 6. Opposite is the case for $Q=2$, which has the lowest caloric effect, with its maximum at a higher temperature, thus being the least advantageous for the caloric phenomenology.
VII Conclusions
This study investigated the magnetocaloric effect (MCE) in a $Q$-state clock model on a square lattice using Monte Carlo simulations, focusing on spin systems with varying $Q$ orientations. By examining $Q$ values from 2 to 8 across different lattice sizes under free boundary conditions, we quantified the MCE through isothermal changes in the external magnetic field, using the system entropy as a key metric.
Our results demonstrate that the magnetocaloric effect can be significantly enhanced through precise control of the spin degrees of freedom. Specifically, for $Q\geq 5$, the system undergoes a double-phase transition characterized by two distinct peaks in specific heat, which correlate with an enhanced caloric response. At lower temperatures ($T<1$ ), systems with $Q=6$ and $Q=8$ exhibit a notable caloric effect driven by the ferromagnetic (FP) to Berezinskii–Kosterlitz–Thouless (BKT) transition. Although this transition occurs at different temperatures for each $Q$ value, the resulting caloric response remains nearly identical, indicating that increasing the spin orientations shifts the transition temperature while preserving the effect’s magnitude.
The behavior observed for $Q=4$ is particularly noteworthy, as it presents a single, pronounced peak in the specific heat, maximizing the MCE at a single critical temperature point. Unlike higher $Q$ systems, which involve more complex double-phase transitions, the $Q=4$ case offers an efficient entropy change with simplified thermal control, suggesting its potential utility in applications where streamlined caloric responses are advantageous.
In summary, this study provides a deeper understanding of how spin orientations within the $Q$-state clock model influence the magnetocaloric effect (MCE), offering critical insights for optimizing caloric responses in magnetic refrigeration and other thermodynamic systems. These findings not only advance the theoretical comprehension of phase transitions in multi-state systems but also reveal practical strategies for harnessing these phenomena across diverse domains of statistical physics such as neural networks, biological systems, and percolation theory. The framework established here opens new pathways for future research into optimizing thermodynamic efficiency in spin-based materials, with promising implications for a range of applied technologies.
VIII Acknowledgments
M. A., F. J. P., E. E. V., and P. V. acknowledge the financial support of ANID Fondecyt (Chile) under contracts 1230055 and 1240582. S. P. A. acknowledges the financial support of UTFSM DPP. F.J.P. acknowledges financial support from “Millennium Nucleus in NanoBioPhysics” project NNBP NCN2021 _021 and UTFSM DGIIE. N. C. acknowledges financial support from ANID Iniciación en Investigación Fondecyt Grant No. 11221088 and DGII-UTA.
Appendix A Exact and Mean Field Approximation
In this section, we analyze two cases where the MCE effect calculation comes from an entropy that can be derived directly from a canonical partition function $(Z)$. The first corresponds to an exact calculation of all the accessible microstates of the system for a $3\times 3$ lattice for $Q=2,4,6$ and $8$. In this framework, it is then possible to obtain $Z$ analytically, and hence, all thermodynamic quantities can be derived from it.